OpenStudy (anonymous):

Vector A (with arrow above) has a negative x component 3.26 units in length and a positive y component 1.94 units in length. (a) Determine an expression for A (with arrow above) in unit-vector notation. (__i+__j) units (b) Determine the magnitude and direction of A(with arrow above). magnitude_____unit(s) direction______° counterclockwise from the positive x axis (c) What vector B(with arrow above) when added to vector A (with arrow above), gives a resultant vector with no x component and a negative y component 4.8 units in length? (__i+__j) units

OpenStudy (anonymous):

a) x component goes before 'i' while y component goes before 'j'

OpenStudy (anonymous):

so it would be xi+yj?

OpenStudy (anonymous):

yeah, but you have to put actual numbers

OpenStudy (anonymous):

or 3.26i+1.94j

OpenStudy (anonymous):

almost it says " negative x component" so it should be -3.26i+1.94j

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to find magnitude , take the square root of sum of square (of x and y) \[\sqrt{(-3.26)^2+\left(1.94\right)^2}\]

OpenStudy (anonymous):

b)magnitude of a=(3sqrt58)/10

OpenStudy (anonymous):

i got 3.79

OpenStudy (anonymous):

direction would be 1.95/3.26?

OpenStudy (anonymous):

yes, magnitude is correct

OpenStudy (anonymous):

arc tan?

OpenStudy (anonymous):

for direction , you take arctan

OpenStudy (anonymous):

\[\text{Arctan}\left(\frac{1.94}{3.26}\right)\] this will get you a terminal angle

OpenStudy (anonymous):

i got 329.24...bc i got a negative and had to add 360

OpenStudy (anonymous):

this is the angle we want 

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we got that angle to be 30.7565